You invited 56 people to your graduation party. You can afford to rent 5 tables, round
and/or rectangular (each costing the same). Each round table can seat 8 people and
each rectangular table can seat 12 people. How many round and rectangular and round
tables should you rent?
1
Expert's answer
2020-03-05T16:46:52-0500

Here, it is given that total number of person invited is 56

and total 5 table can be rented

Now, let x be the number of round table and y be the number of rectangular table

So,

as per the given condition

x+y=5 ......................................1

and for second condition it is given that 8 people can seat around round table and 12 around rectangular table

so, for second condition

8x +12y = 56 ..........................2

from equation 1 , x= 5-y

and substitute this value of x in equation 2

8(5-y) +12 y= 56

40-8y +12y= 56

4y=56-40

4y=16y

so, y= 4

now put this valye of y in 1 we get , x= 1

so total round table be 1 and total rectangular table be 4